Time Remaining 2 hours 55 minutes 19 seconds

02:55:19

Item 1

Item 1

Time Remaining 2 hours 55 minutes 19 seconds

Assume that you manage a risky portfolio with an expected rate of return of 17% and a standard deviation of 27%. The T-bill rate is 7%.

You estimate that a passive portfolio invested to mimic the S&P 500 stock index provides an expected rate of return of 13% with a standard deviation of 25%.

a. What is the slope of the CML?
Answer #1

a). Slope = [E(rM) - rF] /

= [0.13 - 0.07] / 0.25 = 0.06 / 0.25 = 0.24